OpenStudy (anonymous):

What is the equation of the line that contains the point (–5, –1) and has a slope of 4? Write in slope-intercept form. A. y = 4x – 19 B. y = 4x + 19 C. y = 4x – 1 D. y = 4x + 1

OpenStudy (roadjester):

Are you familiar with this: \(\large y-y_0=m(x-x_0)\)?

OpenStudy (anonymous):

yaa

OpenStudy (roadjester):

Ok so you have \((x_0, y_0\)) which is (-5, -1)

OpenStudy (roadjester):

m=4 You just need to substitute.

OpenStudy (roadjester):

Does that make sense?

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

The answer is B. y=4+19

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

No problem.